# Exact analytical algorithm for solvent accessible surface area and derivatives in implicit solvent molecular simulations on GPUs

### v1

@@ Abstract
In this paper, we present dSASA (differentiable SASA), an exact geometric method to calculate solvent accessible surface area (SASA) analytically along with atomic derivatives on GPUs. The atoms in a molecule are first assigned to tetrahedra in groups of four atoms by Delaunay tetrahedrization adapted for efficient GPU implementation and the SASA values for atoms and molecules are calculated based on the tetrahedrization information and inclusion-exclusion method. The SASA values from the numerical icosahedral-based method can be reproduced with more than 98% accuracy for both proteins and RNAs. Having been implemented on GPUs and incorporated into the software Amber, we can apply dSASA to implicit solvent molecular dynamics simulations with inclusion of this nonpolar term. The current GPU version of GB/SA simulations has been accelerated up to nearly 20-fold compared to the CPU version and it outperforms LCPO as the system size increases. The performance and importance of the nonpolar part in implicit solvent modeling are demonstrated in GB/SA simulations of proteins and accurate SASA calculation of nucleic acids.
@@ 1 Introduction
where the polar term $\Delta G_{pol}$ gives the difference between the work of uncharging the solute in vacuum and the work of charging the solute in solvent, $\Delta G_{cav}$ denotes the accommodation of cavity in the solvent for solute and $\Delta G_{vdW}$ represents the van der Waals interactions between solute and solvent. The Poisson-Boltzmann (PB) equation models the electric field for the electrostatic interactions between solvent and solute 8 and can be used to compute $\Delta G_{pol}$. The $\Delta G_{cav}$ and $\Delta G_{vdW}$ are usually combined into one nonpolar term $\Delta G_{np}$. It is considered to be proportional to the number of atoms in the solute having direct contact with solvent molecules and can be estimated in terms of the solvent accessible surface area (SASA) 9. Although it was pointed out that incorporation of the volume term into the computation can provide a more complete description of the nonpolar part 10, 11, the surface area based methods can still have good performance in the prediction of the native-like conformations of proteins and the estimation of ligand-binding affinities 12.
@@
In implicit solvent molecular dynamics (MD) simulations, most of the effort was devoted to the development of the dominant polar part. Due to the complexity of solving PB equation, the generalized Born (GB) method 5 was proposed to approximate PB equation with simple functions. The improvement of GB model on proteins 13, 14, 15 and the variant to nucleic acids 16 and its implementation on graphics processing units (GPUs) 17 helped GB achieve popularity in recent years. The main goal of GB model is to reproduce the “perfect” effective radii and the free energies estimated from PB model. In particular, the pairwise decomposition version of GB 18 is ideal for implementation on GPU with parallel computing. Although the performance of GB is improving, poor folding stabilities were observed in folding studies 2, structure predictions 3 and de novo designed peptides with high helicity 19. As hypothesized in 2, the ignored nonpolar solvation term might result in instability; this was supported by improved fold stability when an approximate SASA-based nonpolar term was added to the same GB model20. As shown in that paper, the inclusion of the nonpolar term in the GB/SA MD simulations can produce more stable trajectories and better simulated melting temperatures than GB-only simulations.
@@
There exist several methods to calculate the molecular surface area on CPUs and GPUs. However, only a few of them can be applied to MD simulations because most of them lack reliable derivatives for atoms. The solvent accessible surface was proposed by Lee and Richards in 21. A numerical implementation by Connolly in 22 is computationally expensive and lacks analytical derivatives. The derivatives are unreliable or unavailable in other methods, such as grid based, neighbor-counting or machine learning 23. The approximation methods on GPU have similar issues 24, 25. In CHARMM software, GBMV2/SA 26 can reproduce the molecular surface and has been implemented on GPUs 27. However, the computation is based on grid points so lower resolution of the grids may generate less reliable results and higher resolution may result in slower simulation speed. The first available method for approximating surface area along with derivatives in Amber is LCPO 28 which is able to estimate SASA based on the neighbors of atoms. As the MD simulations in Amber are now primarily on GPUs 29, 17, the data transfer between CPU and GPU deteriorates the performance because LCPO is implemented only on CPUs. Recently, another approximate pairwise method with a faster speed (up to  30 times faster), called pwSASA, was proposed and it achieved a comparable accuracy to LCPO according to the testing of proteins in 20. The main purpose of pwSASA was to investigate the impact of the nonpolar term on the MD simulations of proteins with a simple and fast pairwise approximation suitable for running entirely on GPUs. The simple form of the functions provided faster simulation speed and did prove the significance of nonpolar term in the simulations. However, the computed SASA values had various correlation coefficients from 0.6 to 0.9 compared with the numerical computation for a number of proteins, which is partly because the higher order interactions were ignored. Besides the varied performance for different protein systems, a parametrization for nucleic acids or small molecule ligands is not available in pwSASA because it was trained exclusively based on proteins.
@@
As every atom in a molecule can be represented by a sphere, the surface areas have also been studied by computational geometry methods. One reliable and accurate method was developed based on the Alpha Complex construction 30 by which the geometric descriptors of molecules can be computed through union of balls. The Alpha Complex is obtained with the atomic coordinates and radii information and can be applied to decompose the space into small segments through which the surface areas are computed by using Gauss-Bonnet theorem 31 or inclusion-exclusion formulas 32, 33, 34. Recently, the extension to other geometric measures, such as volumes, mean and Gaussian curvatures, was proposed in 35 where the parallel implementation can achieve speed improvement over the serial version. However, the full implementation of such method on GPUs is still unavailable because each step in the method requires substantial effort to utilize the parallel computing property to improveh performance.
@@
Here we present dSASA, a new method to calculate SASA and its derivatives with respect to atomic coordinates using Alpha Complex theory and inclusion-exclusion method. This fully analytical and accurate algorithm has been implemented on GPUs and incorporated in Amber software. The basic background and the procedures in the method will be first shown. The results for the assessment of the method are next given: the accuracy of estimation is examined by several protein systems with diverse topologies, the speed of the GPU version in GB/SA simulations is compared with the CPU version and other methods, the performance in GB/SA simulations on proteins by including this nonpolar term and the validation on RNAs. The findings prove the valuable role of an accurate nonpolar term in improving the ability to predict native-like structures using the implicit solvent simulations.
@@ 2.1 Theory and estimations for nonpolar solvation
The relatively accurate SASA values can be calculated by ICOSA numerical method 36 ($gbsa=2$ in Amber), in which starting from an icosahedron a water probe with radius 1.4 Å is recursively rolled on the van der Waals surface of the molecule. The implementation of this method in current Amber MD simulations is unavailable because of the absence of atomic forces. The linear combinations of pairwise overlaps (LCPO; $gbsa=1$ in Amber) is the first stable algorithm used in GB/SA MD simulations in Amber. The neighbor list of a central atom is used to subtract the pairwise overlapping from its isolated sphere areas, which results in the computation complexity $O(N^{2})$, so it will slow down quadratically as the size of the molecule increases. In the GPU version of GB/SA MD with this method, the data transfer between CPU and GPU decelerates the simulations as its implementation is CPU/GPU hybrid. The pairwise approximation of SASA (pwSASA; $gbsa=3$ in Amber) 20 is the recently developed algorithm to calculate SASA that is suitable for implementing on GPUs. The idea is similar to LCPO but only considers the interactions between two atoms, which would achieve a faster speed but result in inaccuracy especially when the molecule contains some unusual atoms. The parameters for the atom types are trained exclusively on proteins so the application to other systems, such as nucleic acids or small molecules, is limited.
@@
Step 2. Create and classify the dual complex $\mathcal{C}$ which denotes the possible interactions among close atoms. The complex $\mathcal{C}$ contains the lists of simplices, such as the vertices (the atoms), edges (interactions between two atoms), triangles (interactions among three atoms) and tetrahedra (interactions among four atoms) from the tetrahedrization. These simplices are then filtered and classified into interior and exterior (denoted by $\partial\mathcal{C}$) based on their connection information.
@@
In the following, we show the process to compute the atomic SASA based on the exterior simplices $\partial\mathcal{C}$ because only the atoms having direct contacts with solvent contribute to the surface areas. The simplices have been classified into interior and exterior at Step 2. The Laguerre diagram of the point set is conjugate to the wDT. The space can be decomposed into cells by the conjugated Laguerre diagram as shown in Figure 1. The Laguerre cell for a weighted point (atom) $p_{i}^{\prime}$ is defined by
@@
Figure 2: Inclusion-exclusion process of $p_{i}$. The red circle denotes the standalone ball $S_{i}$ of $p_{i}$; the shaped region indicates the intersection $S_{ij}$ and $S_{ik}$ with $p_{j},p_{k}$; the small red region is the intersection of three atoms $S_{ijk}$.
@@
By substituting Eqs. (3), (4), (5), (6) and (7) into Eqs. (1) and (2), we can obtain the surface area and the derivative for every atom. More details of the description and the equations can be found in 34, 33 . However, the atomic derivatives may become discontinuous when atoms are approaching. Through detailed analysis in 33 , it can be seen that most common type of discontinuities occur when two atoms become externally tangent. We examine here the simplest case where only two atoms with extended radii $r_{i}=2.9,r_{j}=3.1$ are considered to show the possible singularity in the calculation. The changing process of the SASA values and the atomic derivatives is given in Figure 4. In the left, the summation of SASA values is continuous as two atoms are approaching from a long distance and the minimal value occurs when two atoms are internally tangent. In the right, the magnitude of the atomic derivatives has a leap when two atoms are externally tangent at which the singularity occurs. As investigated in 33 , the singularity rarely happens when the solvent radius is 1.4 Å. In the implementation, we need one threshold value $\epsilon$ to determine the external tangency if $||p_{i}-p_{j}|-(r_{i}+r_{j})|<\epsilon$. When $\epsilon=10^{-6}$, the external tangency was not detected after simulating a few systems for thousands of steps. When $\epsilon=10^{-3}$, it may have one such case after dozens of steps simulation. On the other hand, the magnitude of the leap in the atomic derivatives is less than 20 Å${}^{2}$ and the surface tension is usually less than 0.01 kcal/($\cdot$ Å${}^{2}$), so the impact of such singularity is insignificant compared to the overall forces for exterior atoms whose magnitude can be more than 10 kcal/($\cdot$ Å${}^{2}$). Given the little chance of happening and the tiny impact of the singularity on the overall forces of the atoms, we can expect that the discontinuity in SASA calculation will have negligible impact on the trajectories in GB/SA simulations.
Figure 4: Trend of SASA and the magnitude of atomic derivatives as the distance of two atomic centers is changing. The x-axis shows the distance between the atomic centers. The summation of SASA values is on the left, while the atomic derivatives are on the right.
2.3 Implementation of dSASA on GPU
@@ 2.4 Simulated protein systems, RNA systems
The CD experiments provide the melting curves for Trp-cage 39 and the melting temperature for homeodomain variant 40. These two protein systems were studied in 20 and the ab initio folding experiments in 2 using the same force field and solvent model, providing a good reference to quantify the possible improvement by addition of a nonpolar solvation term.
@@ 2.5 Details of MD simulation
In the GB and GB/SA MD simulations, replica exchange molecular dynamics (REMD) was applied to enhance the efficiency of sampling. The setting for the system is as follows: all bonds involving hydrogen were added SHAKE constraints; Langevin dynamics ($ntt=3$) with 1 $ps^{-1}$ collision frequency was used; the time step was 4 fs by following the protocol in 43 through which the masses of hydrogen atoms can be repartitioned; exchanges between adjacent temperature replicas were attempted every 1 ps; conformations were extracted every 0.1 ns. Trp-cage and homeodomain were parametrized by ff14SBonlysc 44 with GBNeck2 15 and mbondi3 radii. The surface tension parameter was extensively tested in pwSASA 20 for proteins and 5, 7 and 10 cal/($\cdot$ Å${}^{2}$) were working properly in Amber. Then we implemented the GB/SA simulations for LCPO, pwSASA and dSASA with 5 or 7 cal/($\cdot$ Å${}^{2}$).
For Trp-cage, two production runs starting from unfolded and the first NMR structure were simulated for 1.0 $\mu s$ in GB, LCPO, pwSASA and dSASA, with a REMD ladder of 8 temperatures (247.7, 264.0, 281.4, 300.0, 319.8, 340.9, 363.3, 387.3 K) . The backbone RMSD cutoff 2.0 Å was applied to calculate the fraction of folded. For homeodomain variant, two production runs starting from extended and the first NMR structure were simulated with a ladder of 10 temperatures (288.7, 300.0, 311.7, 323.9, 336.6, 349.8, 363.5, 377.7, 392.4, 407.8 K) for 2.0 $\mu s$ in GB and pwSASA, 1.0 $\mu s$ for LCPO and 0.7 $\mu s$ for dSASA. The backbone RMSD cutoff 5.0 Å was used to calculate the fraction of folded. These cutoffs are consistent with the study in 2, 20. MD simulations were not performed for the RNA systems, since weaknesses in current RNA force fields make it challenging to obtain stable simulations for stable hairpins such as 2KOC even with fully explicit water 45.

### v2

@@ Abstract
In this paper, we present dSASA (differentiable SASA), an exact geometric method to calculate solvent accessible surface area (SASA) analytically along with atomic derivatives on GPUs. The atoms in a molecule are first assigned to tetrahedra in groups of four atoms by Delaunay tetrahedrization adapted for efficient GPU implementation and the SASA values for atoms and molecules are calculated based on the tetrahedrization information and inclusion-exclusion method. The SASA values from the numerical icosahedral-based method can be reproduced with more than 98% accuracy for both proteins and RNAs. Having been implemented on GPUs and incorporated into the software Amber, we can apply dSASA to implicit solvent molecular dynamics simulations with inclusion of this nonpolar term. The current GPU version of GB/SA simulations has been accelerated up to nearly 20-fold compared to the CPU version, outperforming LCPO, a commonly used, fast algorithm for calculating SASA, as the system size increases. While we focus on the accuracy of the SASA calculations for proteins and nucleic acids, we also demonstrate stable GB/SA MD mini-protein simulations.
@@ 1 Introduction
where the polar term $\Delta G_{pol}$ gives the difference between the work of uncharging the solute in vacuum and the work of charging the solute in solvent, $\Delta G_{cav}$ denotes the accommodation of cavity in the solvent for solute and $\Delta G_{vdW}$ represents the van der Waals interactions between solute and solvent. The Poisson-Boltzmann (PB) method describes the electrostatic potential in solution for a given set of boundary conditions and determines the effect of electrostatic interactions on the molecules 8 and can be used to compute $\Delta G_{pol}$. The $\Delta G_{cav}$ and $\Delta G_{vdW}$ are usually combined into one nonpolar term $\Delta G_{np}$. It is considered to be proportional to the number of atoms in the solute having direct contact with solvent molecules and can be estimated in terms of the solvent accessible surface area (SASA) 9. Although it has been pointed out that incorporation of the volume term into the computation can provide a more complete description of the nonpolar contribution 10, 11, the surface area based methods can still have good performance in the prediction of the native-like conformations of proteins and the estimation of ligand-binding affinities 12.
@@
In implicit solvent molecular dynamics (MD) simulations, most of the effort has been devoted to the development of the dominant polar part. Due to the complexity of solving the PB equation, the generalized Born (GB) method 5 was proposed to approximate the solution of the PB equation with simple functional forms. The recent improvement of fast GB models for proteins 13, 14, 15 and nucleic acids 16 and its implementation on graphics processing units (GPUs) 17 helped GB achieve popularity in recent years. The main goal of a GB model is to achieve agreement of electrostatic free energies with the PB approach, which can be achieved by parametrizing the GB screening parameters to provide a better match of the effective radii to those from the reference method. In particular, the pairwise decomposition version of GB 18 is ideal for implementation on GPU with parallel computing. Although the performance of GB is improving, poor folding stabilities were observed in protein folding studies 2, structure predictions 3 and de novo designed peptides with high helicity 19. As hypothesized in 2, the neglected nonpolar solvation term might result in instability; this was supported by improved fold stability when an approximate SASA-based nonpolar term was added to the same GB model20. As shown in that paper, the inclusion of the nonpolar term in the GB/SA MD simulations can produce more stable trajectories and better simulated melting temperatures than comparable simulations with identical setup other than the SASA term.
@@
There exist several methods to calculate the molecular surface area on CPUs and GPUs. However, only a few of them can be applied to MD simulations because most of them lack reliable derivatives for atoms. The solvent accessible surface was proposed by Lee and Richards 21. A numerical implementation by Connolly 22 is computationally expensive and lacks analytical derivatives. The derivatives are unreliable or unavailable in other methods, such as grid based, neighbor-counting or machine learning 23. The approximation methods on GPU have similar issues 24, 25. In the CHARMM software, GBMV2/SA 26 can reproduce the molecular surface and has been implemented on GPUs 27. However, the computation is based on grid points so lower resolution of the grids may generate less reliable results while higher resolution results in slower simulation speed. The first available method for approximating surface area along with derivatives in Amber is LCPO 28 which is able to estimate SASA based on the neighbors of atoms. As the MD simulations in Amber are now primarily on GPUs 29, 17, the data transfer between CPU and GPU deteriorates the performance because LCPO is implemented only on CPUs. Recently, another approximate pairwise method with a faster speed (up to  30 times faster), called pwSASA, was proposed and it achieved a comparable accuracy to LCPO according to the testing of proteins 20. The main purpose of pwSASA was to investigate the impact of the nonpolar term on the MD simulations of proteins with a simple and fast pairwise approximation suitable for running entirely on GPUs. The simple form of the functions provided faster simulation speed and did prove the significance of nonpolar term in the simulations. However, the computed SASA values had various correlation coefficients from 0.6 to 0.9 compared with the numerical computation for a number of proteins, which is partly because the higher order interactions were ignored. Furthermore, pwSASA is highly empirical, and the parameters were trained only for the specific local atomic environments founds in amino acids (such as hybridization and the identity of bound atoms). Thus in addition to relatively low accuracy on proteins, pwSASA cannot be used in simulations of more diverse systems such as those including modified amino acids, small molecule ligands, nucleic acids, or glycans, further limiting its suitability.
@@
As every atom in a molecule can be represented by a sphere, the surface areas have also been studied by computational geometry methods. One reliable and accurate method was developed based on the Alpha Complex construction 30 by which the geometric descriptors of molecules can be computed through union of balls. The Alpha Complex is obtained with the atomic coordinates and radii information and can be applied to decompose the space into small segments through which the surface areas are computed by using Gauss-Bonnet theorem 31 or inclusion-exclusion formulas 32, 33, 34. Recently, the extension to other geometric measures, such as volumes, mean and Gaussian curvatures, was proposed 35 in which the parallel implementation can achieve speed improvement over the serial version. However, the full implementation of such method on GPUs is still unavailable because each step in the method requires substantial effort to utilize the parallel computing property to improve performance.
@@
Here we present dSASA, a new method to calculate SASA and its derivatives with respect to atomic coordinates using Alpha Complex theory and inclusion-exclusion method. This fully analytical and accurate algorithm has been implemented on GPUs and incorporated in Amber software. The basic background and the procedures employed in the method will be first shown. The results for the assessment of the method are next given: the accuracy of estimation is examined by several protein systems with diverse topologies, the speed of the GPU version in GB/SA simulations is compared with the CPU version and other methods, and the performance is demonstrated by including this nonpolar term in GB/SA simulations on proteins. However, we caution readers that good practice in force field development involves isolation and validation of the individual components as done here. Deviations from experimental behavior can arise from inaccuracy in any of the multiple components (solute force field, polar solvation term, and nonpolar term). Likewise, a good match to experiment can be obtained from fortuitous cancellation of error. For this reason, we show that the dSASA term performs well in MD simulations, which verifies the efficient algorithm implementation, but our main focus remains on validating the SASA accuracy.
@@ 2.1 Theory and estimations for nonpolar solvation
The relatively accurate SASA values can be calculated by ICOSA numerical method 36 ($gbsa=2$ in Amber), in which starting from an icosahedron a water probe with radius 1.4 Å is recursively rolled on the van der Waals surface of the molecule. The implementation of this method in current Amber MD simulations is unavailable because of the absence of atomic forces. The Linear Combinations of Pairwise Overlaps (LCPO; $gbsa=1$ in Amber) is the first stable algorithm used in GB/SA MD simulations in Amber. The neighbor list of a central atom is used to subtract the pairwise overlapping from its isolated sphere areas, which results in the computation complexity $O(N^{2})$, where $N$ is the total number of solute atoms, so it will slow down quadratically as the size of the molecule increases. In the GPU version of GB/SA MD with this method, the data transfer between CPU and GPU decelerates the simulations as its implementation is CPU/GPU hybrid. The pairwise approximation of SASA (pwSASA; $gbsa=3$ in Amber) 20 is the recently developed algorithm to calculate SASA that is suitable for implementing on GPUs. The idea is similar to LCPO but only considers the interactions between two atoms, which achieves a faster speed but results in an approximate SASA. The parameters for the atom types are trained exclusively on standaard amino acids, so the application to other systems, such as nucleic acids or small molecules, is unavailable.
@@
Step 2. Create and classify the dual complex $\mathcal{C}$ which denotes the possible interactions among close atoms. The complex $\mathcal{C}$ contains the lists of simplices, such as the vertices (the atoms), edges (overlap between two atoms), triangles (overlap among three atoms) and tetrahedra (overlap among four atoms) from the tetrahedrization. These simplices are then filtered and classified into interior and exterior (denoted by $\partial\mathcal{C}$) based on their connection information.
@@
In the following, we show the process to compute the atomic SASA based on the exterior simplices $\partial\mathcal{C}$ because only the atoms having direct contact with solvent contribute to the surface areas. The simplices have been classified into interior and exterior at Step 2. The Laguerre diagram of the point set is conjugate to the wDT. The space can be decomposed into cells by the conjugated Laguerre diagram as shown in Figure 1. The Laguerre cell for a weighted point (atom) $p_{i}^{\prime}$ is defined by
@@
Figure 2: Inclusion-exclusion process of $p_{i}$. The red circle denotes the standalone ball $S_{i}$ of $p_{i}$; the shaded region indicates the intersection $S_{ij}$ and $S_{ik}$ with $p_{j},p_{k}$; the small red region is the intersection of three atoms $S_{ijk}$.
@@
By substituting Eqs. (3), (4), (5), (6) and (7) into Eqs. (1) and (2), we can obtain the surface area and the derivative for every atom. More details of the description and the equations can be found in 34, 33 . However, the atomic derivatives may become discontinuous when atoms are approaching. Through detailed analysis in 33, it can be seen that most common type of discontinuities occur when two atoms become externally tangent. We examine here the simplest case where only two atoms with extended radii $r_{i}=2.9,r_{j}=3.1$ are considered to show the possible singularity in the calculation. The changing process of the SASA values and the atomic derivatives is given in Figure 4. In the left, the summation of SASA values is continuous as two atoms are approaching from a long distance and the minimal value occurs when two atoms are internally tangent. In the right, the magnitude of the atomic derivatives has a leap when two atoms are externally tangent at which the singularity occurs. As investigated in 33, the singularity rarely happens when the solvent radius is 1.4 Å. In the implementation, we need one threshold value $\epsilon$ to determine the external tangency if $||p_{i}-p_{j}|-(r_{i}+r_{j})|<\epsilon$. When $\epsilon=10^{-6}$, the external tangency was not detected after simulating a few systems for thousands of steps. When $\epsilon=10^{-3}$, there may occur one such case after dozens of simulation steps. On the other hand, the magnitude of the leap in the atomic derivatives is less than 20 Å2 and the surface tension is usually less than 0.01 kcal/($\cdot$ Å2), so the impact of such singularity is insignificant compared to the overall forces for exterior atoms whose magnitude can be more than 10 kcal/mol. Given the little chance of happening and the tiny impact of the singularity on the overall forces of the atoms, we can expect that the discontinuity in SASA calculation will have negligible impact on the trajectories in GB/SA simulations.
Figure 4: Trend of SASA and the magnitude of atomic derivatives as the distance of two atomic centers is changing. The x-axis shows the distance between the atomic centers, decreasing to the right. The summation of SASA values is on the left, while the atomic derivatives are on the right.
2.3 Implementation of dSASA on GPUs
@@ 2.4 Simulated protein systems, RNA systems
The CD experiments provide the melting curves for Trp-cage 39 and the melting temperature for homeodomain variant 40. These two protein systems were studied in 20 and the ab initio folding experiments in 2 using the same force field and solvent model, providing a good reference to quantify the possible improvement by addition of a nonpolar solvation term. As mentioned above, it is important to compare results in which only the component of interest is varied.
@@ 2.5 Details of MD simulation
In the GB and GB/SA MD simulations, replica exchange molecular dynamics (REMD) was applied to enhance the efficiency of sampling. The setting for the system is as follows: all bonds involving hydrogen were added SHAKE constraints; Langevin dynamics ($ntt=3$) with 1 $ps^{-1}$ collision frequency was used; the time step was 4 fs by following the protocol in 43 through which the masses of hydrogen atoms can be repartitioned; exchanges between adjacent temperature replicas were attempted every 1 ps; conformations were extracted every 0.1 ns. Trp-cage and homeodomain were parametrized by ff14SBonlysc 44 with GBNeck2 15 and mbondi3 radii 15. The surface tension parameter was extensively tested in pwSASA 20 for proteins and 5, 7 and 10 cal/($\cdot$ Å2) provided reasonable results when compared to simulations of a model system in explicit water. Thus we implemented the GB/SA simulations for LCPO, pwSASA and dSASA with 5 or 7 cal/($\cdot$ Å2).
For Trp-cage, two production runs starting from unfolded and the first NMR structure were simulated for 1.0 $\mu s$ in GB, LCPO, pwSASA and dSASA, with a REMD ladder of 8 temperatures (247.7, 264.0, 281.4, 300.0, 319.8, 340.9, 363.3, 387.3 K) . The backbone RMSD cutoff 2.0 Å was applied to calculate the fraction of folded. For homeodomain variant, two production runs starting from extended and the first NMR structure were simulated with a ladder of 10 temperatures (288.7, 300.0, 311.7, 323.9, 336.6, 349.8, 363.5, 377.7, 392.4, 407.8 K) for 2.0 $\mu s$ in GB and pwSASA, 1.0 $\mu s$ for LCPO and dSASA. The backbone RMSD cutoff 5.0 Å was used to calculate the fraction of folded. These cutoffs are consistent with the study in 2, 20. MD simulations were not performed for the RNA systems, since weaknesses in current RNA force fields make it challenging to obtain stable simulations for stable hairpins such as 2KOC even with fully explicit water 45. We therefore restrict our analysis to the SASA accuracy for RNA.